# EMOVA[Uncaptioned image]: Empowering Language Models to See, Hear and Speak with Vivid Emotions

### v2

@@ 6 Conclusion
This work presents EMOVA, an innovative end-to-end omni-modal large language model that effectively aligns vision, speech, and text simultaneously. We employ a continuous vision encoder to capture fine-grained visual details, while a discrete, semantic-acoustic disentangled speech tokenizer and detokenizer enable end-to-end speech understanding and generation. A lightweight style module further supports spoken dialogue with vivid emotions. By using text as a bridge, we demonstrate that omni-modal alignment is achievable without relying on scarce omni-modal image-text-speech data, which not only enhances both vision-language and speech capabilities but also surpasses its bi-modal counterparts through joint optimization. For the first time, EMOVA achieves state-of-the-art performance on both vision-language and speech benchmarks, setting a novel standard for the omni-modal models for versatile and expressive omni-modal interactions.
@@ A.1 Speech-to-Unit (S2U) Tokenizer
To process the speech input $\textbf{X}_{S}$, our S2U tokenizer consists of a speech encoder $s(\cdot)$ with a quantization module $q(\cdot)$. First, the speech input is passed through $s(\cdot)$, producing a continuous latent representation $\textbf{E}_{S}=s(\textbf{X}_{S})$. Then, the quantization module $q(\cdot)$ converts $\textbf{E}_{s}$ into discrete units $\textbf{U}_{S}=q(\textbf{E}_{s})$. The final output is an ID sequence $\textbf{U}_{S}=[u_{1},u_{2},\cdots]$, where each $u_{i}$ corresponds to a unique speech unit.
After this S2U extraction process, the speech is represented by quantized IDs instead of quantized embeddings. For example, a speech signal is represented as an ID sequence like $[782,463,550,\cdots]$, which can be treated as a special form of text. As a result, to integrate speech tokens into LLM $f(\cdot)$, we only need to expand the LLM’s original vocabulary $V_{T}$ by adding a set of speech unit tokens $V_{S}$, similar to Zhang et al. (2023a). The expanded vocabulary is thus the union $V=V_{T}\cup V_{S}$. In this work, the same codebook is shared across multiple languages, such as English and Chinese, enabling the unit modeling abilities to be shared across languages. This design enables simply integration of the speech modality into the LLM with minimal data requirements (see experiments in Sec. 4.1).
@@
The training of the S2U encoder involves three primary components: the speech encoder, the quantization module, and the phone decoder. First, the speech encoder is trained with a large amount of unlabeled speech with contrastive loss in a self-supervised learning manner (Tao et al., 2024). The dataset utilized is the 10000-hour English speeches from GigaSpeech (Chen et al., 2021a) and the 10000-hour Chinese speeches from WenetSpeech (Zhang et al., 2022), both of which encode large variations in speakers, styles, and acoustic environments. Subsequently, the speech encoder, along with the quantization module and phone decoder, are optimized on a small amount of speech-text pair data, which is derived from the train-clean-100 subset of LibriSpeech (Panayotov et al., 2015) in English and AISHELL-1 (Bu et al., 2017) in Chinese. The phone label is obtained using the phone set in Charsiu (Zhu et al., 2022). During training, the speech encoder encodes input speeches into continuous latent representations that are rich in both phonetic and tonal information. Then, the quantization module is used to convert the continuous outputs from the speech encoder into discrete units. Finally, the phone decoder converts the quantized embeddings into a sequence of non-tonal/tonal phones, ensuring that the speech units capture necessary information related to semantic contents in both non-tonal and tonal languages. After training, only the speech encoder and the quantization module are used in EMOVA.
@@
Our S2U tokenizer differs from the SpeechTokenizer (Zhang et al., 2023b) used in AnyGPT (Zhan et al., 2024), in the following aspects:
@@
(2) Compared with SpeechTokenizer, our S2U tokenizer offers a more concise representation and helps to simplify and accelerate the generation of EMOVA. SpeechTokenizer employs tokens from eight RVQ layers with a 50Hz frame rate to represent speech, thus a 10-second speech corresponds to 500 × 8 = 4000 tokens. However, we reduce the frame rate from 50Hz to 25Hz and utilize only one token to represent each frame, and thus, a 10-second speech can be represented by only 250 tokens. Moreover, AnyGPT requires a two-stage generation process, involving autoregressive (AR) semantic token generation followed by the non-autoregressive (NAR) acoustic token generation. Instead, we only need to generate speech units capturing semantic contents in a fully AR manner.
@@ A.2 Unit-to-Speech (U2S) Detokenizer with Style Control
The proposed U2S detokenizer involves three core modules: the speech unit encoder $e(\cdot)$, the speech style encoder $g(\cdot)$, and the speech decoder $d(\cdot,\cdot)$. First, the speech unit encoder converts the speech units $\textbf{U}_{S}^{o}$ into unit embeddings $\textbf{E}_{unit}^{o}$. Meanwhile, the style encoder $g(\cdot)$, adopting the structure of Meta-StyleSpeech (Min et al., 2021), is utilized to extract a style embedding $\textbf{E}_{style}^{o}$ from the chosen reference speech. Lastly, the speech decoder $d(\cdot,\cdot)$ reconstructs the speech waveform $\textbf{Y}_{S}^{o}$ from the unit embedding $\textbf{E}_{unit}^{o}$ and style embedding $\textbf{E}_{style}^{o}$.
@@
Training a U2S detokenizer with emotion controls is challenging considering the lack of labeled emotional speech data since most open-source speech data is predominantly neutral in emotion or lacks emotion labels. Due to the limited availability of emotion-rich data, we utilize TTS tools (Du et al., 2024) to generate a small set of style-rich speech samples diverse in speaker identities, genders, emotions, and pitches. Our U2S detokenizer is first pre-trained on LibriTTS (Zen et al., 2019) and AISHELL-1 (Bu et al., 2017) to acquire fundamental speech synthesis capabilities, and subsequently, the synthesized style-rich speech data is utilized to fine-tune the U2S detokenizer, enhancing its controllability over diverse speech styles.
@@
To better facilitate controls of genders, emotions, and pitches, inspired by Min et al. (2021) that a small number of style reference speeches can effectively transfer the target styles, we adopt a “store-for-usage” manner, i.e., we construct a style prototype codebook in advance for speech style assignation. Specifically, we synthesize $K$ reference candidates with external TTS tools for each possible combination of the following styles: two genders (male, female), seven emotions (neutral, happy, sad, angry), and three pitches (normal, high, low), leading to 24 unique styles and $24\times K$ candidates. Empirically we find that genders and pitches are easy to control using any of the candidate references, while the emotion intensity varies across speeches. To tackle this, we adopt Emotion2Vec (Ma et al., 2023), a powerful speech emotion recognition (SER) tool, to measure the emotion intensity of each candidate reference, and rank them in terms of the confidence of the desired emotion. We select the Top-1 candidate reference in each combination style to be the prototype of this condition. Finally, the most representative 24 reference speeches are selected from the $24\times K$ candidates.
@@ B.1 Omni-modal Text-Centric Alignment
The training configuration adopted in Sec. 4.1 is mostly identical to Table except that we use a unique resolution of 448 for all stages and replace EMOVA-SFT in Stage-3 with ShareGPT4V (Chen et al., 2023d) for efficiency.
@@ B.2 Omni-modal Instruction Data Synthesis
As discussed in Sec. 4.2, the omni-modal instruction data is formulated as $D_{\text{omni}}=\{(x_{V},u_{S},x^{o}_{T},c^{o}_{\text{style}},u^{o}_{S})_{i}\}_{i=1}^{N}$. In details, the textual outputs $x_{T}^{o}=(x_{T}^{o^{1}},x_{T}^{o^{1}})$ contain the transcribed textual instructions $x_{T}^{o^{1}}$ and the textual responses $x_{T}^{o^{2}}$. The styles labels $c^{o}_{\text{style}}=(c^{o}_{\text{emo}},c^{o}_{\text{p}})$ include the emotion and pitch labels, respectively.
@@
Based on the notations above, the sequential chain of modality approach can be mathematically formulated by decomposing the conditional likelihood of the desired outputs $(x_{T}^{o^{1}},x_{T}^{o^{1}},c^{o}_{\text{emo}},c^{o}_{\text{p}},u^{o}_{S})$ given the inputs $(x_{V},u_{S})$. Specifically, let $z_{1}=x_{T}^{o^{1}},z_{2}=x_{T}^{o^{1}},z_{3}=c^{o}_{\text{emo}},z_{4}=c^{o}_{\text{p}},\text{and},z_{5}=u^{o}_{S}$, the decomposition is expressed as:
$\mathbb{P}(x_{T}^{o^{1}},x_{T}^{o^{1}},c^{o}_{\text{emo}},c^{o}_{\text{p}},u^{o}_{S}\mid x_{V},u_{S})=\prod_{i=1}^{5}\mathbb{P}(z_{i}\mid z_{1:i-1},x_{V},u_{S}).$ (2)
Table 5: Statistics of the EMOVA speech instruction tuning datasets.
@@
In-house 686 96 196 198 196 458 134 92
@@ Appendix C More on Benchmark Evaluation
To thoroughly evaluate our model’s vision-language abilities, 14 benchmarks covering four different aspects of real-life scenarios are utilized for a comprehensive assessment across multiple domains. Moreover, Automatic Speech Recognition (ASR) and Text-to-speech (TTS) are adopted to evaluate speech-language abilities.
@@
Benchmarks including the TextVQA (Singh et al., 2019), ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022), and OCRBench (Liu et al., 2023c), focus on recognition and understanding of structured data (e.g., charts, documents, and characters), challenging the model to extract, comprehend, and reason with structural and textual data. We adopt the corresponding test set for evaluation.
@@
MME (Fu et al., 2024a), MMBench (Liu et al., 2023b), SEED-Bench-Image (Li et al., 2023a), MM-Vet (Yu et al., 2024), and RealWorldQA (xAI, 2024) evaluate models on perception and reasoning among general vision domains, providing a comprehensive measurement of models’ real-world generalization abilities.
@@
MathVista (testmini) (Lu et al., 2024) and MathVerse (vision-mini) (Zhang et al., 2024) evaluate the model’s ability to deal with diverse mathematical problems, including both arithmetic and complex reasoning questions across different levels of complexity.
@@
ScienceQA-Img (Lu et al., 2022) and AI2D (Kembhavi et al., 2016) are utilized to assess the model’s ability to deal with scientific questions and diagrams, which requires reasoning across various subjects and interpreting structured visual content.
@@
We utilize the test-clean set of LibriSpeech (Panayotov et al., 2015) for English, reporting the Word Error Rate (WER) as the evaluation metric. For Chinese, evaluation is conducted on the test set of AISHELL-2 (Du et al., 2018), using the Character Error Rate (CER). Both WER and CER assess ASR performance, calculated by comparing the recognized texts with the ground-truth transcripts.
@@
To evaluate the TTS abilities, we first prompt EMOVA to generate speech units, which are then converted to speech waveforms by the U2S detokenizer. Using the synthesized speech as input, we conduct ASR with Whisper-large-v3 and Paraformer-zh for English and Chinese, respectively, to obtain transcribed texts. We then compute the WER and CER between the ground truth texts and the transcribed texts as metrics for TTS. The resulting metrics are denoted as TTS-WER and TTS-CER for English and Chines.
@@ D.1 Calculation of Evaluation Metrics
of our U2S detokenizer is assessed by providing texts to LLM to generate corresponding units (i.e., TTS), which, along with the given conditional emotion labels, are then fed into our U2S detokenizer to synthesize speech. We choose female voice due to its large variation of styles. We evaluate on 4 commonly-seen emotion, i.e.,“neutral”, “happy”, “sad”, and “angry”. We synthesize 200 speech utterances for testing, with 50 utterances per emotion. The output speeches are analyzed by a Speech Emotion Recognition (SER) model named Emotion2Vec (Ma et al., 2023), which identifies the emotion with the greatest likelihood among these four emotion.
@@ D.2 Comparison with other Omni Models
Experimental results of ASR and TTS are reported in Table 3 and 6. EMOVA achieves significant improvements over other omni-modal models (i.e., AnyGPT and VITA), even surpassing its SLLM counterpart Mini-Omni (Xie & Wu, 2024), demonstrating the effectiveness of semantic-acoustic disentanglement and omni-modal mutual benefits. For the first time, our EMOVA obtains state-of-the-art performance on both the vision-language and speech benchmarks simultaneously.
Appendix E More on Vision-language

### v3

@@ 6 Conclusion
Our work builds EMOVA, a novel end-to-end omni-modal large language model that effectively aligns vision, speech, and text simultaneously. With text as a bridge, we show that omni-modal alignment is achievable without relying on omni-modal image-text-speech data, meanwhile, enhancing both vision-language and speech abilities. For the first time, EMOVA achieves state-of-the-art performance on both vision-language and speech benchmarks, setting a new standard for versatile omni-modal interactions.
Acknowledgments.
We gratefully acknowledge supports of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research. This research has been made possible by funding support from the Research Grants Council of Hong Kong through the Research Impact Fund project R6003-21.
@@ A.1 Speech-to-Unit (S2U) Tokenizer
To process the speech input $\textbf{X}_{S}$, our S2U tokenizer consists of a speech encoder $s(\cdot)$ with a quantization module $q(\cdot)$. First, the speech input is passed through $s(\cdot)$, producing a continuous latent representation $\textbf{E}_{S}=s(\textbf{X}_{S})$. Then, the quantization module $q(\cdot)$ converts $\textbf{E}_{s}$ into discrete units $\textbf{U}_{S}=q(\textbf{E}_{s})$. The final output is an ID sequence $\textbf{U}_{S}=[u_{1},u_{2},\cdots]$, where each $u_{i}$ corresponds to a unique speech unit in the speech codebook.
After this S2U extraction procedure, the speech is represented by quantized IDs instead of quantized embeddings. For example, a speech signal is represented as an ID sequence like $[782,463,550,\cdots]$, which can be treated as a special form of text. As a result, to integrate speech tokens into LLM $f(\cdot)$, we only need to expand the LLM’s original vocabulary $V_{T}$ by adding a set of speech unit tokens $V_{S}$, similar to [94]. The expanded vocabulary is thus the union $V=V_{T}\cup V_{S}$. In this work, the same codebook is shared across multiple languages, such as English and Chinese, enabling the unit modeling abilities to be shared across languages. This design enables simply integration of the speech modality into the LLM with minimal data requirements (see experiments in Sec. 4.1).
@@
The training of the S2U encoder involves three primary components: the speech encoder, the quantization module, and the phone decoder. First, the speech encoder is trained with a large amount of unlabeled speech with contrastive loss in a self-supervised learning manner [77]. The dataset utilized is the 10000-hour English speeches from GigaSpeech [3] and the 10000-hour Chinese speeches from the WenetSpeech [93], both of which encode large variations in speakers, styles, and acoustic environments. Subsequently, the speech encoder, along with the quantization module and phone decoder, are optimized on a small amount of speech-text pair data, which is derived from the train-clean-100 subset of LibriSpeech [72] in English and AISHELL-1 [2] in Chinese. The phone label is obtained using the phone set in Charsiu [98]. During training, the speech encoder encodes input speeches into continuous latent representations that are rich in both phonetic and tonal information. Then, the quantization module is used to convert the continuous outputs from the speech encoder into discrete units. Finally, the phone decoder converts the quantized embeddings into a sequence of non-tonal/tonal phones, ensuring that the speech units capture necessary information related to semantic contents in both non-tonal and tonal languages. After training, only the speech encoder and the quantization module are used in EMOVA.
@@
Our S2U tokenizer differs from the SpeechTokenizer [96] adopted in AnyGPT [92], in the following aspects:
@@
(2) Compared with SpeechTokenizer, our S2U tokenizer offers a more concise representation and helps to simplify and accelerate generation. SpeechTokenizer employs tokens from eight RVQ layers with a 50Hz frame rate to represent speech, thus a 10-second speech corresponds to 500 × 8 = 4000 tokens. However, we reduce the frame rate from 50Hz to 25Hz and utilize only one token to represent each frame, and thus, a 10-second speech can be represented by only 250 tokens. Moreover, AnyGPT requires a two-stage generation process, involving autoregressive (AR) semantic token generation followed by the non-autoregressive (NAR) acoustic token generation. Instead, we only need to generate speech units capturing the semantic contents in a fully AR manner.
@@ A.2 Unit-to-Speech (U2S) Detokenizer with Style Control
The proposed U2S detokenizer involves three core modules: the speech unit encoder $e(\cdot)$, the speech style encoder $g(\cdot)$, and the speech decoder $d(\cdot,\cdot)$. First, the speech unit encoder converts the speech units $\textbf{U}_{S}^{o}$ into unit embeddings $\textbf{E}_{unit}^{o}$. Meanwhile, the style encoder $g(\cdot)$, adopting the structure of Meta-StyleSpeech [66], is utilized to extract a style embedding $\textbf{E}_{style}^{o}$ from the chosen reference speech. Lastly, the speech decoder $d(\cdot,\cdot)$ reconstructs the speech waveform $\textbf{Y}_{S}^{o}$ from the unit embedding $\textbf{E}_{unit}^{o}$ and style embedding $\textbf{E}_{style}^{o}$.
@@
Training a U2S detokenizer with emotion controls is challenging considering the lack of labeled emotional speech data since most open-source speech data is predominantly neutral in emotion or lacks emotion labels. Due to the limited availability of emotion-rich data, we utilize TTS tools [16] to generate a small set of style-rich speech samples diverse in speaker identities, genders, emotions, and pitches. Our U2S detokenizer is first pre-trained on LibriTTS [91] and AISHELL-1 [2] to acquire fundamental speech synthesis capabilities, and subsequently, the synthesized style-rich speech data is utilized to fine-tune the U2S detokenizer, enhancing its controllability over diverse speech styles.
@@
To better facilitate controls of genders, emotions, and pitches, inspired by [66] that a small number of style reference speeches can effectively transfer the target styles, we adopt a “store-for-usage" manner, i.e., we construct a style prototype codebook in advance for speech style assignation. Specifically, we synthesize $K$ reference candidates with external TTS tools for each possible combination of the following styles: two genders (male, female), four emotions (neutral, happy, sad, angry), and three pitches (normal, high, low), leading to 24 unique styles and $24\times K$ candidates. Empirically we find that genders and pitches are easy to control using any of the candidate references, while the emotion intensity varies across speeches. To tackle this, we adopt Emotion2Vec [59], a powerful speech emotion recognition (SER) tool, to measure the emotion intensity of each candidate reference, and rank them in terms of the confidence of the desired emotion. We select the Top-1 candidate reference in each combination style to be the prototype of this condition. Finally, the most representative 24 reference speeches are selected from the $24\times K$ candidates.
@@ B.1 Omni-modal Text-Centric Alignment
The training configuration adopted in Sec. 4.1 is mostly identical to Table 5 except we use a unique resolution of 448 for all stages and replace EMOVA-SFT-4M in Stage-3 with ShareGPT4V [9] for efficiency.
@@ B.2 Omni-modal Instruction Data Synthesis
As previously discussed in Sec. 4.2, the omni-modal instruction data is formulated as $D_{\text{omni}}=\{(x_{V},u_{S},x^{o}_{T},c^{o}_{\text{style}},u^{o}_{S})_{i}\}_{i=1}^{N}$. In details, the textual outputs $x_{T}^{o}=(x_{T}^{o^{1}},x_{T}^{o^{2}})$ contain the transcribed textual instructions $x_{T}^{o^{1}}$ and the textual responses $x_{T}^{o^{2}}$. The styles labels $c^{o}_{\text{style}}=(c^{o}_{\text{emo}},c^{o}_{\text{p}})$ include the emotion and pitch labels, respectively.
@@
Based on the notations above, the sequential chain of modality approach can be mathematically formulated by decomposing the conditional likelihood of the desired outputs $(x_{T}^{o^{1}},x_{T}^{o^{2}},c^{o}_{\text{emo}},c^{o}_{\text{p}},u^{o}_{S})$ given the inputs $(x_{V},u_{S})$. Specifically, let $z_{1}=x_{T}^{o^{1}},z_{2}=x_{T}^{o^{2}},z_{3}=c^{o}_{\text{emo}},z_{4}=c^{o}_{\text{p}},\text{and},z_{5}=u^{o}_{S}$, the decomposition is expressed as:
$\mathbb{P}(x_{T}^{o^{1}},x_{T}^{o^{2}},c^{o}_{\text{emo}},c^{o}_{\text{p}},u^{o}_{S}\mid x_{V},u_{S})=\prod_{i=1}^{5}\mathbb{P}(z_{i}\mid z_{1:i-1},x_{V},u_{S}).$ (3)
@@
In-house 684 96 196 196 196 458 134 92
Table 4: Statistics of the EMOVA speech instruction tuning datasets.
Settings Stage-1 Stage-2 Stage-3
Vision Resolution Min $56^{2}$, Max $1792^{2}$ Min $56^{2}$, Max $1792^{2}$ Min $56^{2}$, Max $1792^{2}$
# Tokens Min $4$, Max $4096$ Min $4$, Max $4096$ Min $4$, Max $4096$
Data Dataset LCS EMOVA-Alignment (Fig. 9) EMOVA-SFT (Fig. 8)
# Samples 558K 7.4M 4.4M
Training Trainable Projector Full Model (Half ViT) Full Model
Batch Size 256 256 128
Context Length 8192 8192 8192
LR: $p(\cdot)$ $1\times 10^{-3}$ $2\times 10^{-5}$ $2\times 10^{-5}$
LR: $v(\cdot)$ $2\times 10^{-5}$ $2\times 10^{-6}$
LR: $f(\cdot)$ $2\times 10^{-5}$ $2\times 10^{-5}$
Epoch 1 1 1
@@ Appendix C More on Benchmark Evaluation
To thoroughly evaluate our model’s vision-language abilities, 14 benchmarks covering four different aspects of the real-life scenarios are utilized for a comprehensive assessment across multiple domains. Moreover, Automatic Speech Recognition (ASR) and Text-to-speech (TTS) are adopted to evaluate speech-language abilities.
@@
Benchmarks including TextVQA [76], ChartQA [61], DocVQA [62], InfoVQA [63], and OCRBench [51], focus on recognition and understanding of structured data (e.g., charts, documents, and characters), challenging the model to extract, comprehend, and reason with structural and textual data. We adopt the corresponding test set for evaluation.
@@
MME [18], MMBench [50], SEED-Bench-Image [39], MM-Vet [90], and RealWorldQA [86] evaluate models on perception and reasoning among general vision domains, providing a comprehensive measurement of models’ real-world generalization abilities.
@@
MathVista (testmini) [57] and MathVerse (vision-mini) [95] evaluate the model’s ability to deal with diverse mathematical problems, including both arithmetic and complex reasoning questions across different levels of complexity.
@@
ScienceQA-Img [56] and AI2D [33] are utilized to assess the model’s ability to deal with scientific questions and diagrams, which requires reasoning across various subjects and interpreting structured visual content.
@@
We utilize the test-clean set of LibriSpeech [72] for English, reporting the Word Error Rate (WER) as the evaluation metric. For Chinese, evaluation is conducted on the test set of AISHELL-2 [15], using the Character Error Rate (CER). Both WER and CER assess ASR performance, calculated by comparing the recognized texts with the ground-truth transcripts.
@@
To evaluate the TTS abilities, we first prompt EMOVA to generate speech units, which are then converted to the speech waveforms by our U2S detokenizer. Using the synthesized speech as input, we conduct ASR with Whisper-large-v3 and Paraformer-zh for English and Chinese, respectively, to obtain transcribed texts. We then compute the WER and CER between the ground truth texts and the transcribed texts as metrics for TTS. The resulting metrics are denoted as TTS-WER and TTS-CER for English and Chinese.
@@ D.1 Calculation of Evaluation Metrics
of our U2S detokenizer is assessed by providing texts to LLM to generate corresponding units (i.e., TTS), which, along with the given conditional emotion labels, are then fed into our U2S detokenizer to synthesize speech. We choose the female voice due to its large variation of styles. We evaluate on 4 commonly-seen emotions, i.e.,“neutral”, “happy”, “sad”, and “angry”. We synthesize 200 speech utterances for testing, with 50 utterances per emotion. The output speeches are analyzed by a Speech Emotion Recognition (SER) model named Emotion2Vec [59], which identifies the emotion with the greatest likelihood among these four emotions.
[…]
@@ D.2 Comparison with other Omni Models
Experimental results of ASR and TTS are reported in Table 2 and 6. EMOVA achieves significant improvements over other omni-modal models (i.e., AnyGPT and VITA), even surpassing its SLLM counterpart Mini-Omni [87], demonstrating the effectiveness of semantic-acoustic disentanglement and omni-modal mutual benefits. For the first time, our EMOVA obtains state-of-the-art performance on both the vision-language and speech benchmarks simultaneously.
Appendix E More on Vision-language Architecture